PLEEEASE HEELP!! PLEASE! Show in two ways that the points (6,1) (2,-3) and (4,-5) are the vertices of a right triangle.
Answer by ewatrrr(24785)


Hi
(6,1)
(2,-3) m = 4/4 = 1 D = sqrt(32), D^2 = 32
(4,-5) m = 2/-2 = -1 Lines perpendicular, right angle, d = sqrt(8), D^2 = 8
(6,1) D = sqrt(36 + 4) = sqrt(40), D^2 = 40
32 + 8 = 40